it takes hank twice as long to build one table as it takes to build one chair. If it takes 30 hours to make a dining room set of 1 table and 4 chairs. how long does it take to make the table?
Answer by oberobic(2304)

t = time to build one table
c = time to build one chair
.
t = 2c
.
t + 4c = 30 hr
.
substitute t = 2c
.
2c +4c = 30
6c = 30
c = 5
.
So the time to build a chair = 5 hr to build one chair
.
t = 2c
t = 2(5)
t = 10 hr to build a table
.
Done.